OpenStudy (anonymous):

A set of equations is given below: Equation A: e = 4f + 1 Equation B: e = 3f + 5 Which of the following is a step that can be used to find the solution to the set of equations? 4f + 1= 3f 4f = 3f + 5 4f + 5 = 3f + 1 4f + 1= 3f + 5

Mehek (mehek14):

any ideas?

OpenStudy (anonymous):

4f+5=3f+1

Mehek (mehek14):

the value of e in the first equation is 4f + 1 so all you need to do is put 4f + 1 for e in the second equation

Mehek (mehek14):

How'd you get 4f + 5?

OpenStudy (anonymous):

is it the first one

Mehek (mehek14):

close

Mehek (mehek14):

what happened to +5 after 3f?

Mehek (mehek14):

you keep 4f + 1 as it is and 3f + 5 as it is

OpenStudy (anonymous):

4f + 1= 3f + 5

OpenStudy (anonymous):

The fourth one. 4f+1 = 3f+5